how do i solve this? (-8000)^4/3.This is what i had tried but am now stuck. (#=square root) sorry i dont know the proper abv for it. 3^#8000^4 i know the answer is 160 000 but i dont understand how they got to that answer i tryed everything i can think of. i assume that bc i changed it to that format the negative just goes away. your help will be greatly apreciated (sorry about the spelling doing an all nighter.
Answer by scott8148(5896)

when you have a fractional exponent, the numerator is a power and the denominator is a root

8000^(4/3)=[cuberoot(8000)]^4

cuberoot(8000)=[cuberoot(8)]=2*10=20

20^4=(2^4)*(10^4)=16*10000=160000